PlayQueen:

Find the expression that represents the length of the hypotenuse of a right triangle whose legs measure 2004-04-01-02-00_files/i0350000.jpgand 2004-04-01-02-00_files/i0350001.jpg. A. 2004-04-01-02-00_files/i0350002.jpg B. 2004-04-01-02-00_files/i0350003.jpg C. 2004-04-01-02-00_files/i0350004.jpg D. 2004-04-01-02-00_files/i0350005.jpg

Aries:

Do you think you can add the pictures on here?

Aries:

On this

PlayQueen:

1 attachment
Aries:

Yes thank you

PlayQueen:

oh ok

Aries:

Pythagorean Theorem a2+b2=c2 a and b are the legs, so we just substitute that into the equation (2x2-y2)2+(2xy)2=c2 First simplify the legs (a and b) =(2x2−y2)2 =4x4−4x2y2+y4 =(2xy)2 =4x2y2 So we now have (4x4−4x2y2+y4)+4x2y2 Combine like terms 4x4−4x2y2+y4+4x2y2 =4x4+y4 So we have 4x4+y4=c2 Now we take the square root on both sides c=4x4+y4−−−−−−−√ Which happens to be A

Aries:

Does that make sense?